Consider the following reaction: CaCO3(s)→CaO(s)+CO2(g). Estimate ΔG∘ for this reaction at each of the following temperatures. (Assume that ΔH∘ and ΔS∘ do not change too much within the given temperature range.)

A) 320 K

B) 1045 K

C) 1435 K

Predict whether each reaction will be spontaneous or not.
Answer #1

- | cao, (9 Cao (8) + Co, (4) 14° = JAH (ca 013] + Aug [CO, (9) } - Amj [caco, co] = (-634.8+(-3835)-(-120716) AH = 1492 kJ